OpenStudy (anonymous):

The equation y=-1777x+27153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is x=0 corresponds to 2000,x=2 corresponds to 2002 x=5 corresponds to 2005 and so on. predict the number of gun deaths in 2005 and 2007. In what year will the number of gun deaths be 12937?

OpenStudy (saifoo.khan):

y=-1777x+27153 2005 y =-1777(5)+27153 y = 18,268 2007 y =-1777(7)+27153 y = 14,714

OpenStudy (saifoo.khan):

in 2008, the gun deaths will be 12937

OpenStudy (pokemon23):

saifoo do my question I like your work ;0

OpenStudy (saifoo.khan):

lol, Alright.. i didnt wanted to intrupt.